OpenStudy (anonymous):

Simplify: 18 – 4 + 12 ÷ 2

OpenStudy (ragingsquirrel):

Do you know the order of operations?

OpenStudy (anonymous):

1 13 15 20 are the only choices though

OpenStudy (mertsj):

Whoops sorry. Thought the divide sign was a positive.

OpenStudy (mertsj):

Do the division first and the work in order from left to right.

OpenStudy (mertsj):

18-4+6=14+6=20

OpenStudy (ragingsquirrel):

The order of operations can be abbreviated PEMDAs. I didn't think it would help me a while ago, but it did. Parentheses Equations Multiplication Division Addition

OpenStudy (ragingsquirrel):

Just handle those in order.

OpenStudy (ragingsquirrel):

Whoops, i mean exponents.

OpenStudy (anonymous):

Answer is last option, 20.